# Perpetual Demand Lending Pools

### v1

@@ Abstract
Decentralized perpetuals protocols have collectively reached billions of dollars of daily trading volume, yet are still not serious competitors on the basis of trading volume with centralized venues such as Binance. One of the main reasons for this is the high cost of capital for market makers and sophisticated traders in decentralized settings. Recently, numerous decentralized finance protocols have been used to improve borrowing costs for perpetual futures traders. We formalize this class of mechanisms utilized by protocols such as Jupiter, Hyperliquid, and GMX, which we term Perpetual Demand Lending Pools (PDLPs). We then formalize a general target weight mechanism that generalizes what GMX and Jupiter are using in practice. We explicitly describe pool arbitrage and expected payoffs for arbitrageurs and liquidity providers within these mechanisms. Using this framework, we show that under general conditions, PDLPs are easy to delta hedge, partially explaining the proliferation of live hedged PDLP strategies. Our results suggest directions to improve capital efficiency in PDLPs via dynamic parametrization.
@@ 1 Introduction
Perpetual futures markets are the most liquid trading markets for cryptocurrencies. These markets facilitate daily volumes on the order of hundreds of billions of dollars of notional value and are the primary drivers of hedging and leverage. Although the majority of perpetual futures volume remains on centralized trading venues, such as Binance or Coinbase, a growing amount of perpetuals trading occurs on decentralized venues. Such venues, including Hyperliquid, Jupiter, GMX, and dYdX, have grown to capture over 10% of the daily perpetual future volume [12].
@@
PDLP LPs have had success delta heading in practice. These LPs, who provide assets to the pool, can take derivative positions to offset their exposure to risky asset price movement (the delta) while still capturing fees. This property allows LPs to enter and exit positions without realizing large convexity losses due to volatility [35] and, as a result, facilitates cheaper market making [33, 55]. We note that many previous attempts in decentralized finance to provide delta-hedged portfolios have been unsuccessful in practice (e.g., see [5, 42, 45]). This observation begs the question: Why are PDLPs easier to hedge? We characterize PDLP mechanics and shed light on this question in this paper.
@@
Finally, in §4, we describe dynamic delta hedging of risk assets in a PDLP using mean-variance optimization. We give simple sufficient conditions for when the delta hedge improves the Sharpe ratio of PDLP LP returns. We also describe when a single PDLP pool should be split into two pools to improve the delta hedged Sharpe ratio.
@@ 2 Perpetual Demand Lending Pools
In this section, we formalize a model of Perpetual Demand Lending Pools (PDLPs). Note that we ignore implementation details of practical mechanisms that do not impact economic payoffs.. We begin by describing a simplified linear model of a perpetuals exchange. We refer the interested reader to [6] for models that account for additional parameters found in practice.
@@ 2.1 Perpetuals exchange
In words, if the the loss in the perpetuals position due to price change is greater than the collateral value, the position will be liquidated. For example, a long position is liquidated once the price of the risk asset drops below $p_{0}-c/\delta$. If this position was created with minimal collateral, saturating (1), then the position is liquidated when the price drops below $p_{0}(1-1/\eta)$.
@@
Consider a trade $T=(c,\delta,\eta,p_{0})$ with collateral $c=\$2000$, position size $\delta=1$ ETH, leverage $\eta=4$ and initial ETH price $p_{0}=\$2000$. The initial notional value of the position, $\delta p_{0}$, is 8,000 USD. This trade satisfies the collateral condition (1). As the price of ETH $p$ decreases, the position remains valid until liquidation condition is met, which occurs when $p=\$1,5000$. Once this price is readched, the perpetuals exchange will liquidate this position.
@@
Consider an ETH-USD perpetual futures contract $(L,S,p_{0})$ with cumulative long position $L=1000$ ETH, short position $S=250$ ETH, and mark price $p_{0}=2000$ USD, the funding rate is $\gamma_{L}=\kappa\left(\frac{L}{S}-\frac{p}{p_{0}}\right)$. When $p=p_{0}$, including at the start of the period, this equation simplifies to:
@@ 2.2 Perpetual Demand Lending Pools
Formally, we specify a Perpetual Demand Lending Pool (PDLP) by its current portfolio $R\in{\mbox{\bf R}}^{n}_{+}$, target portfolio $\pi\in{\mbox{\bf R}}^{b}_{+}$, lending fee $f\in(0,1)$, and outstanding loans $\{c_{i}\}$, where each loan $c_{i}$ is the collateral used by position $i=1,\dots,M$. Note that one can view the loans made by the PDLP as collateralizing an equal but opposite position to the one desired by the trader, much as a market marker takes the opposite size of a leveraged trade. The trader associated with position $i$ must pay a fee $fc_{i}$ for each time period that they hold a loan. We note that the target portfolio and fees may be dynamically updated based on external market conditions. The total value lent to traders must be less than the total value of the pool’s assets, i.e.,
@@ Discussion.
The same price movement also introduces a discrepancy between the price at which the PDLP is willing to exchange assets and the external market price. An arbitrageur may buy the asset at the old price $p_{0}$ (which need not be the mark price) on the PDLP and sell it on an an external market for the new, higher price $p$. The price impact of this trade on both markets and the associated fees bound the size of this trade. For simplicity, we focus only on the price impact of buying on the PDLP. (We will return to the PDLP’s fees in §3.) We model the PDLP’s swap functionality using the forward exchange function $G:{\mbox{\bf R}}_{+}\to{\mbox{\bf R}}_{+}$, which is nonnegative, concave, and nondecreasing [1]. This function denotes, for a quantity $x$ of the numéraire, the amount of the asset, $G(x)$ that the PDLP is willing to sell. We assume that $G(0)=0$ (no free lunch) and that the PDLP initially quotes at the old price: $G^{\prime}(0)=1/p_{0}$. If the arbitraguer uses $x$ units of the numéraire to purchase the asset on the PDLP, their profit is
@@ 3.1 Target Weight Mechanisms
minimize $\left\|w(R+\Delta)-w^{\star}\right\|,$ (6)
@@
with variable $\Delta\in{\mbox{\bf R}}^{n}$ denoting the assets to be added or removed from the PDLP. The constraint indicates that the amount removed cannot be in excess of the unutilized assets $R^{A}$. The act of providing incentives to approximately solve (6) resembles both the Hedge Algorithm [25] and other online learning algorithms on the simplex [32]. We suspect that weight update rules could be designed to be no-regret, as they are for in resource markets in [4, 18]. We leave formalizing this connection to future work.
@@
Here, we walk through a simple example to demonstrate the benefits of the target weight mechanism. Consider a PDLP with two assets, the numéraire and a risky asset, that are equally weighetd. Assume that all of the risky asset is being used as collateral for perpetual positions. What happens when the price of the risky asset changes? If the risky asset decreases in price, the PDLP will be underweight the risk asset. Arbitrageurs will be incentivitized to add more of this asset to the pool, which will allow for more positions that use the risky asset as collateral to be opened. On the other hand, if the risky asset increases in price, the PDLP will be overweight the risky asset. Arbitrageurs can either close positions in the money, redeeming the risky asset, or add numéraire to the pool.
@@ 3.2 TWMs in practice
GMX GLP pool and Jupiter JLP pool
. GMX introduced the first PDLP, launched on the Arbitrum blockchain, in September 2021 [56]. This PDLP provided incentives for users to maintain a target weight: changes to the pool that brought the asset composition closer to the target received a rebate, whereas changes that pushed the asset composition further from the target incurred a fee. As implemented, the mechanism resembles a simple proportional-integral-derivative (PID) controller mechanism that requires arbitrageurs to implement the control policy [58]. Jupiter’s JLP pool introduced a similar mechanism but with changes to reduce the funding rate risk borne by LPs using delta-hedge strategies. As of December 2024, GMX’s PDLP has approximately \$650 million in assets [19] and supports \$10 million in open interest (dollar notional value of open positions). Jupiter’s PDLP is significantly larger, with roughly \$1.5 billion of assets and \$870 million of open interest [27].
@@ Appendix A Trading Dynamics
Appendix B When is it better to have a multiple pools?

### v2

@@ Abstract
Decentralized perpetuals protocols have collectively reached billions of dollars of daily trading volume, yet are still not serious competitors on the basis of trading volume with centralized venues such as Binance. One of the main reasons for this is the high cost of capital for market makers and sophisticated traders in decentralized settings. Recently, numerous decentralized finance protocols have been used to improve borrowing costs for perpetual futures traders. These protocols have grown to over \$2.5 billion dollars of assets while generating over \$890 million in fees in 2024. We formalize this class of mechanisms utilized by protocols such as Jupiter, Hyperliquid, and GMX, which we term Perpetual Demand Lending Pools (PDLPs). We then formalize a general target weight mechanism that generalizes what GMX and Jupiter are using in practice. We explicitly describe pool arbitrage and expected payoffs for arbitrageurs and liquidity providers within these mechanisms. Using this framework, we show that under general conditions, PDLPs are easy to delta hedge, partially explaining the proliferation of live hedged PDLP strategies. Our results suggest directions to improve capital efficiency in PDLPs via dynamic parametrization.
@@ 1 Introduction
Perpetual futures markets are the most liquid trading markets for cryptocurrencies. These markets facilitate daily volumes on the order of hundreds of billions of dollars of notional value and are the primary drivers of hedging and leverage. Although the majority of perpetual futures volume remains on centralized trading venues, such as Binance or Coinbase, a growing amount of perpetuals trading occurs on decentralized venues. Such venues, including Hyperliquid, Jupiter, GMX, and dYdX, have grown to capture almost 10% of the daily perpetual future volume [12].
@@
PDLP LPs have had success delta hedging in practice. These LPs, who provide assets to the pool, can take derivative positions to offset their exposure to risky asset price movement (the delta) while still capturing fees. This property allows LPs to enter and exit positions without realizing large convexity losses due to volatility [35] and, as a result, facilitates cheaper market making [33, 55]. We note that many previous attempts in decentralized finance to provide delta-hedged portfolios have been unsuccessful in practice (e.g., see [5, 42, 45]). This observation begs the question: Why are PDLPs easier to hedge? We characterize PDLP mechanics and shed light on this question in this paper.
@@
Finally, in §4, we describe dynamic delta hedging of risk assets in a PDLP using mean-variance optimization. We give simple sufficient conditions for when the delta hedge improves the Sharpe ratio of PDLP LP returns. We also describe when a single PDLP pool should be split into two pools to improve the delta-hedged Sharpe ratio.
@@ 2 Perpetual Demand Lending Pools
In this section, we formalize a model of Perpetual Demand Lending Pools (PDLPs). Note that we ignore implementation details of practical mechanisms that do not impact economic payoffs. We begin by describing a simplified linear model of a perpetuals exchange. We refer the interested reader to [6] for models that account for additional parameters found in practice.
@@ 2.1 Perpetuals exchange
In words, if the the loss in the perpetuals position due to price change is greater than the collateral value, the position will be liquidated. For example, a long position is liquidated once the price of the risky asset drops below $p_{0}-c/\delta$. If this position was created with minimal collateral, saturating (1), then the position is liquidated when the price drops below $p_{0}(1-1/\eta)$.
@@
Consider a trade $T=(c,\delta,\eta,p_{0})$ with collateral $c=\$2000$, position size $\delta=1$ ETH, leverage $\eta=4$ and initial ETH price $p_{0}=\$2000$. The initial notional value of the position, $\delta p_{0}$, is 8,000 USD. This trade satisfies the collateral condition (1). As the price of ETH $p$ decreases, the position remains valid until the liquidation condition is met, which occurs when $p=\$1,5000$. Once this price is reached, the perpetuals exchange will liquidate this position.
@@
Consider an ETH-USD perpetual futures contract $(L,S,p_{0})$ with cumulative long position $L=1000$ ETH, short position $S=250$ ETH, and mark price $p_{0}=2000$ USD, the funding rate is $\gamma_{L}=\kappa\left(\frac{L}{S}-\frac{p}{p_{0}}\right)$. When $p=p_{0}$, including at the start of the contract period, this equation simplifies to:
@@ 2.2 Perpetual Demand Lending Pools
Formally, we specify a Perpetual Demand Lending Pool (PDLP) by its current portfolio $R\in{\mbox{\bf R}}^{n}_{+}$, target portfolio $\pi\in{\mbox{\bf R}}^{b}_{+}$, lending fee $f\in(0,1)$, and outstanding loans $\{c_{i}\}_{i\in[M]}$, where each loan $c_{i}$ is the collateral used by position $i=1,\dots,M$. Note that one can view the loans made by the PDLP as collateralizing an equal but opposite position to the one desired by the trader, much as a market maker takes the opposite size of a leveraged trade. The trader associated with position $i$ must pay a fee $fc_{i}$ for each time period that they hold a loan. We note that the target portfolio and fees may be dynamically updated based on external market conditions. The total value lent to traders must be less than the total value of the pool’s assets, i.e.,
@@ Discussion.
The same price movement also introduces a discrepancy between the price at which the PDLP is willing to exchange assets and the external market price. An arbitrageur may buy the asset at the old price $p_{0}$ (which need not be the mark price) on the PDLP and sell it on an external market for the new, higher price $p$. The price impact of this trade on both markets and the associated fees bound the size of this trade. For simplicity, we focus only on the price impact of buying on the PDLP. (We will return to the PDLP’s fees in §3.) We model the PDLP’s swap functionality using the forward exchange function $G:{\mbox{\bf R}}_{+}\to{\mbox{\bf R}}_{+}$, which is nonnegative, concave, and nondecreasing [1]. This function denotes, for a quantity $x$ of the numéraire, the amount of the asset, $G(x)$ that the PDLP is willing to sell. We assume that $G(0)=0$ (no free lunch) and that the PDLP initially quotes at the old price: $G^{\prime}(0)=1/p_{0}$. If the arbitraguer uses $x$ units of the numéraire to purchase the asset on the PDLP, their profit is
@@ 3.1 Target Weight Mechanisms
minimize $\left\|w(p,R+\Delta)-w^{\star}\right\|,$ (6)
@@
with variable $\Delta\in{\mbox{\bf R}}^{n}$ denoting the assets to be added or removed from the PDLP. The constraint indicates that the amount removed cannot be in excess of the unutilized assets $R^{A}$. The act of providing incentives to approximately solve (6) resembles both the Hedge Algorithm [25] and other online learning algorithms on the simplex [32]. We suspect that weight update rules could be designed to be no-regret, as they are in resource markets in [4, 18]. We leave formalizing this connection to future work.
@@
Here, we walk through a simple example to demonstrate the benefits of the target weight mechanism. Consider a PDLP with two assets, the numéraire and a risky asset, that are equally weighetd. Assume that all of the risky asset is being used as collateral for perpetual positions. What happens when the price of the risky asset changes? If the risky asset decreases in price, the PDLP will be underweight the risky asset. Arbitrageurs will be incentivitized to add more of this asset to the pool, which will allow for more positions that use the risky asset as collateral to be opened. On the other hand, if the risky asset increases in price, the PDLP will be overweight the risky asset. Arbitrageurs can either close positions in the money, redeeming the risky asset, or add numéraire to the pool.
@@ 3.2 TWMs in practice
GMX GLP pool and Jupiter JLP pool.
GMX introduced the first PDLP, launched on the Arbitrum blockchain, in September 2021 [56]. This PDLP provided incentives for users to maintain a target weight: changes to the pool that brought the asset composition closer to the target received a rebate, whereas changes that pushed the asset composition further from the target incurred a fee. As implemented, the mechanism resembles a simple proportional-integral-derivative (PID) controller mechanism that requires arbitrageurs to implement the control policy [58]. Jupiter’s JLP pool introduced a similar mechanism but with changes to reduce the funding rate risk borne by LPs using delta-hedge strategies. As of December 2024, GMX’s PDLP has approximately \$650 million in assets [19] and supports \$10 million in open interest (dollar notional value of open positions). Jupiter’s PDLP is significantly larger, with roughly \$1.5 billion of assets and \$870 million of open interest [27].
@@ Appendix A Trading Dynamics
Appendix B When is it better to have multiple pools?